1.2g of Silver was roasted in air to form silver 1.289g of silver oxide. Find the Empirical formula for the silver oxide formed.
1
Expert's answer
2012-12-11T08:44:02-0500
Chemical reaction for this process is:

Ag+ x/2 O2= AgxO

If mass of silver = 1.2 and 1.289 is mass of silver oxide, molecular weight of silver oxide is (108x+16)

1.2 & 1.289
Ag+ x/2 O2= AgOx
108 & (108x+16)

108*1.289=1.2*(108+16x)
116.01=108+16x
x=0,5


AgO(0.5) or Ag2O is empirical formula